# Stochastic Gradient Langevin Unlearning

### v1

@@ 3.3 Unlearning Multiple Data Points
Figure 2: Illustration of (a) sequential unlearning and (b) batch unlearning. In both cases, the key idea is to establish an upper bound on the initial $W_{\infty}$ distance. (a) For sequential unlearning, the initial $W_{\infty}$ distance bound $Z_{\mathcal{B}}^{(s)}$ for each $s^{th}$ unlearning request can be derived with triangle inequality. (b) For batch unlearning, we analyze the case that two learning processes can differ in $S\geq 1$ points (Corollary 3.12).
So far we have focused on one unlearning request and unlearning one point. In practice, multiple unlearning requests can arrive sequentially (sequential unlearning) and each unlearning request may require unlearning multiple points (batch unlearning). Below we demonstrate that our stochastic gradient Langevin unlearning naturally supports sequential and batch unlearning as well.
@@
Under the same assumptions as in Corollary 3.8. Assume the unlearning requests arrive sequentially such that our dataset changes from $\mathcal{D}=\mathcal{D}_{0}\rightarrow\mathcal{D}_{1}\rightarrow\ldots\rightarrow\mathcal{D}_{S}$, where $\mathcal{D}_{s},\mathcal{D}_{s+1}$ are adjacent. Let $y_{k}^{j,(s)}$ be the unlearned parameters for the $s^{th}$ unlearning request at $k^{th}$ unlearning epoch and $j^{th}$ iteration following Algorithm (1) on $\mathcal{D}_{s}$ and $y_{0}^{0,(s+1)}=y_{K_{s}}^{0,(s)}\sim\bar{\nu}_{\mathcal{D}_{s}|\mathcal{B}}$, where $y_{0}^{0,(0)}=x_{\infty}$ and $K_{s}$ is the unlearning steps for the $s^{th}$ unlearning request. For any $s\in[S]$, we have $W_{\infty}(\bar{\nu}_{\mathcal{D}_{s-1}|\mathcal{B}},\nu_{\mathcal{D}_{s}|\mathcal{B}})\leq Z_{\mathcal{B}}^{(s)},$ where $Z_{\mathcal{B}}^{(s+1)}=\min(c^{K_{s}n/b}Z_{\mathcal{B}}^{(s)}+Z_{\mathcal{B}},2R)$, $\;Z_{\mathcal{B}}^{(1)}=Z_{\mathcal{B}}$, $Z_{\mathcal{B}}$ is described in Corollary 3.8 and $\;c=1-\eta m$.
@@
By combining Corollary 3.8 and Theorem 3.11, we can establish the least unlearning iterations of each unlearning request $\{K_{s}\}_{s=1}^{S}$ to achieve $(\alpha,\varepsilon)$-RU simultaneously. Notably, our sequential unlearning bound is much better than the one in Chien et al. (2024), especially when the number of unlearning requests is large. The key difference is that Chien et al. (2024) have to leverage weak triangle inequality for Rényi divergence, which double the Rényi divergence order $\alpha$ for each sequential unlearning request. In contrast, since our analysis only requires tracking the initial $W_{\infty}$ distance, where the standard triangle inequality can be applied. As a result, our analysis can better handle the sequential unlearning case. We also demonstrate in Section 4 that the benefit offered by our results is significant in practice.
@@
Batch unlearning. We can extend Lemma 3.5 to the case that adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$ can differ in $S\geq 1$ points. This leads to the following corollary.
Corollary 3.12 ($W_{\infty}$ bound for batch unlearning).
$W_{\infty}(\nu_{T|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$
$\leq\min\left(\frac{1-c^{T}}{1-c}\sum_{g=0}^{G-1}c^{n/b-j_{g}-1}\frac{2\eta MS_{j_{g}}}{b},2R\right).$
@@
Figure 3: Main experiments, where the top and bottom rows are for MNIST and CIFAR10 respectively. (a) Compare to baseline for unlearning one point using limited $K$ unlearning epoch. For SGLU, we use only $K=1$ unlearning epoch. For D2D, we allow it to use $K=1,5$ unlearning epochs. (b) Unlearning $100$ points sequentially versus baseline. For LU, since their unlearning complexity only stays in a reasonable range when combined with batch unlearning of size $S$ sufficiently large, we report such a result only. (c,d) Noise-accuracy-complexity trade-off of SGLU for unlearning $100$ points sequentially with various mini-batch sizes $b$, where all methods achieve $(\epsilon,1/n)$-unlearning guarantee with $\epsilon=0.01$. We also report the required accumulated epochs for retraining for each $b$.
@@ 4 Experiments
Benchmark datasets. We consider binary logistic regression with $\ell_{2}$ regularization. We conduct experiments on MNIST (Deng, 2012) and CIFAR10 (Krizhevsky et al., 2009), which contain 11,982 and 10,000 training instances respectively. We follow the setting of (Guo et al., 2020; Chien et al., 2024) to distinguish digits $3$ and $8$ for MNIST so that the problem is a binary classification. For the CIFAR10 dataset, we distinguish labels $3$ (cat) and $8$ (ship) and leverage the last layer of the public ResNet18 (He et al., 2016) embedding as the data features, which follows the setting of Guo et al. (2020) with public feature extractor.
@@
Baseline methods. Our baseline methods include Delete-to-Descent (D2D) (Neel et al., 2021) and Langevin Unlearning (LU) (Chien et al., 2024), which are the state-of-the-art full-batch gradient-based approximate unlearning methods. Note that when our Stochastic Gradient Langevin Unlearning (SGLU) chooses $b=n$ (i.e., full batch), the learning and unlearning iterations become PNGD which is identical to LU. Nevertheless, the corresponding privacy bound is still different as we leverage the analysis of PABI instead of those based on Langevin dynamics in Chien et al. (2024). Hence, we still treat these two methods differently in our experiment. For D2D, we leverage Theorem 9 and 28 in (Neel et al., 2021) for privacy accounting depending on whether we allow D2D to have an internal non-private state. Note that allowing an internal non-private state provides a weaker notion of privacy guarantee (Neel et al., 2021) and both SGLU and LU by default do not require it. We include those theorems for D2D and a detailed explanation of its possible non-privacy internal state in Appendix L. For LU, we leverage their Theorem 3.2, and 3.3 for privacy accounting (Chien et al., 2024), which are included in Appendix M.
@@
All experimental details can be found in Appendix K, including how to convert $(\alpha,\varepsilon)$-RU to the standard $(\epsilon,\delta)$-unlearning guarantee. Throughout this section, we choose $\delta=1/n$ for each dataset and require all tested unlearning approaches to achieve $(\epsilon,\delta)$-unlearning with different $\epsilon$. We report test accuracy for all experiments as the utility metric. We set the learning iteration $T=10,20,50,1000$ to ensure SGLU converges for mini-batch size $b=32,128,512,n$ respectively. All results are averaged over $100$ independent trials with standard deviation reported as shades in figures.
@@
Unlearning one data point with $K=1$ epoch. We first consider the setting of unlearning one data point using only one unlearning epoch (Figure 3a). For both LU and SGLU, we use only $K=1$ unlearning epoch. Since D2D cannot achieve a privacy guarantee with only limited (i.e., less than $10$) unlearning epoch without a non-private internal state, we allow D2D to have it and set $K=1,5$ in this experiment. Even in this case, SGLU still outperforms D2D in both utility and unlearning complexity. Compared to LU, our mini-batch setting either outperforms or is on par with it. Interestingly, we find that LU gives a better privacy bound compared to full-batch SGLU ($b=n$) and thus achieves better utility under the same privacy constraint, see Appendix M for the detailed comparisons of the privacy bounds. Nevertheless, due to the use of weak triangle inequality in LU analysis, we will see that our SGLU can outperform LU significantly for multiple unlearning requests.
@@
Unlearning multiple data points. Let us consider the case of multiple ($100$) unlearning requests (Figure 3b). We let all methods achieve the same $(1,1/n)$-unlearning guarantee for a fair comparison. We do not allow D2D to have an internal non-private state anymore in this experiment for a fair comparison. Since the privacy bound of LU only gives reasonable unlearning complexity with a limited number of sequential unlearning updates (Chien et al., 2024), we allow it to unlearn $S=10$ points at once. We observe that SGLU requires roughly $10\%$ and $2\%$ of unlearning epochs compared to D2D and LU for $b=n$ and $b=128$ respectively, where all methods exhibit similar utility ($0.9$ and $0.98$ for MNIST and CIFAR10 respectively). It shows that SGLU is much more efficient compared to D2D and LU. Notably, while both SGLU with $b=n$ and LU (un)learn with PNGD iterations, the resulting privacy bound based on our PABI-based analysis is superior to the one pertaining to Langevin-dynamic-based analysis in Chien et al. (2024). See our discussion in Section 3.3 for the full details.
@@
Privacy-utility-complexity trade-off. We now investigate the inherent utility-complexity trade-off regarding noise standard deviation $\sigma$ and mini-batch size $b$ for SGLU under the same privacy constraint, where we require all methods to achieve $(0.01,1/n)$-unlearning guarantee for $100$ sequential unlearning requests (Figure 3c and 3d). We can see that smaller $\sigma$ leads to a better utility, yet more unlearning epochs are needed for SGLU to achieve $\epsilon=0.01$. On the other hand, smaller mini-batch size $b$ requires fewer unlearning epochs $K$ as shown in Figure 3d, since more unlearning iterations are performed per epoch. Nevertheless, we remark that choosing $b$ too small may lead to degradation of model utility or instability. Decreasing the mini-batch size $b$ from $32$ to $1$ reduces the average accuracy of training from scratch from $0.87$ to $0.64$ and $0.97$ to $0.81$ on MNIST and CIFAR10 respectively for $\sigma=0.03$. In practice, one should choose a moderate mini-batch size $b$ to ensure both good model utility and unlearning complexity. Finally, we also note that SGLU achieves a similar utility while much better complexity compared to retraining from scratch, where SGLU requires at most $1,5$ unlearning epochs per unlearning request for $b=32,512$ respectively.
Acknowledgements
@@
The authors thank Sinho Chewi and Wei-Ning Chen for the helpful discussions. E. Chien, H. Wang and P. Li are supported by NSF awards OAC-2117997 and JPMC faculty award.
@@ Appendix A Additional related works
Differential privacy of noisy gradient methods. DP-SGD (Abadi et al., 2016) is arguably the most popular method for ensuring a DP guarantee for machine learning models. Since it leverages the DP composition theorem and thus the privacy loss will diverge for infinite training epochs. Recently, researchers have found that if we only release the last step of the trained model, then we can do much better than applying the composition theorem. A pioneer work (Ganesh & Talwar, 2020) studied the DP properties of Langevin Monte Carlo methods. Yet, they do not propose to use noisy GD for general machine learning problems. A recent line of work (Ye & Shokri, 2022; Ryffel et al., 2022) shows that PNSGD training can not only provide DP guarantees, but also the privacy loss is at most a finite value even if we train with an infinite number of iterations. The main analysis therein is based on the analysis of Langevin Monte Carlo (Vempala & Wibisono, 2019; Chewi, Sinho, 2023). In the meanwhile, (Altschuler & Talwar, 2022a) also provided the DP guarantees for PNSGD training but with analysis based on PABI (Feldman et al., 2018). None of these works study how PNSGD can also be leveraged for machine unlearning.
Appendix B Standard definitions
Let $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ be a mapping. We define smoothness, Lipschitzsness, and strong convexity as follows:
@@ Appendix B Standard definitions
Appendix C Existence of limiting distribution
@@
Suppose that the closed convex set $\mathcal{C}_{R}\subset\mathbb{R}^{d}$ is bounded with $\mathcal{C}_{R}$ having a positive Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}_{R}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. The Markov chain $\{x_{t}:=x_{t}^{0}\}$ in Algorithm 1 for any fixed mini-batch sequence =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=1}^{n/b-1}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on the Borel $\sigma$-algebra of $\mathcal{C}_{R}$. Furthermore, for any $x\in\mathcal{C}_{R}$, the distribution of $x_{t}$ conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$.
The proof is almost identical to the proof of Theorem 3.1 in Chien et al. (2024) and we include it for completeness. We start by proving that the process $\{x_{t}:=x_{t}^{0}\}$ admits a unique invariant measure (Proposition C.1) and then show that the process converges to such measure which is in fact a probability measure (Theorem C.2). Combining these two results completes the proof of Theorem 3.1.
Proposition C.1.
Suppose that the closed convex set $\mathcal{C}_{R}\subset\mathbb{R}^{d}$ is bounded with $\textrm{Leb}(\mathcal{C}_{R})>0$ where Leb denotes the Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}_{R}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. Then the Markov chain $\{x_{t}:=x_{t}^{0}\}$ defined in Algorithm 1 for any fixed mini-batch sequence =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=1}^{n/b-1}$ admits a unique invariant measure (up to constant multiples) on $\mathcal{B}(\mathcal{C}_{R})$ that is the Borel $\sigma$-algebra of $\mathcal{C}_{R}$.
@@
This proposition is a direct application of results from (Meyn & Tweedie, 2012). According to Proposition 10.4.2 in (Meyn & Tweedie, 2012), it suffices to verify that $\{x_{t}\}$ is recurrent and strongly aperiodic.
@@
    Recurrency. Thanks to the Gaussian noise $W_{t}$, $\{x_{t}\}$ is Leb-irreducible, i.e., it holds for any $x\in\mathcal{C}_{R}$ and any $A\in\mathcal{B}(\mathcal{C}_{R})$ with $\textrm{Leb}(A)>0$ that
@@
    where $\tau_{A}=\inf\{t\geq 0:x_{t}\in A\}$ is the stopping time. Therefore, there exists a Borel probability measure $\psi$ such that that $\{x_{t}\}$ is $\psi$-irreducible and $\psi$ is maximal in the sense of Proposition 4.2.2 in (Meyn & Tweedie, 2012). Consider any $A\in\mathcal{B}(\mathcal{C}_{R})$ with $\psi(A)>0$. Since $\{x_{t}\}$ is $\psi$-irreducible, one has $L(x,A)=\mathbb{P}(\tau_{A}<+\infty\ |\ x_{0}=x)>0$ for all $x\in\mathcal{C}_{R}$. This implies that there exists $T\geq 0$, $\delta>0$, and $B\in\mathcal{B}(\mathcal{C}_{R})$ with $\textrm{Leb}(B)>0$, such that $\mathbb{P}(x_{T}\in A\ |\ x_{0}=x)\geq\delta,\ \forall~{}x\in B$. Therefore, one can conclude for any $x\in\mathcal{C}_{R}$ that
@@
    $\geq$ $\sum_{t=1}^{\infty}\delta\cdot\inf_{y\in\mathcal{C}_{R}}\mathbb{P}(x_{t}\in B\ |\ x_{t-1}=y)$
@@
    where we used the fact that $\inf_{y\in\mathcal{C}_{R}}\mathbb{P}(x_{t}\in B\ |\ x_{t-1}=y)=\inf_{y\in\mathcal{C}_{R}}\mathbb{P}(x_{1}\in B\ |\ x_{0}=y)>0$ that is implies by $\textrm{Leb}(B)>0$ and the boundedness of $\mathcal{C}_{R}$ and $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C}_{R};\mathbf{d}_{i})$. Let us remark that we actually have compact $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C}_{R};\mathbf{d}_{i})$ since $\mathcal{C}_{R}$ is compact and $\nabla f(\cdot;\mathbf{d}_{i})$ is continuous. The arguments above verify that $\{x_{t}\}$ is recurrent (see Section 8.2.3 in (Meyn & Tweedie, 2012) for definition).

### v2

@@ 3.2 Unlearning Multiple Data Points
So far we have focused on one unlearning request and unlearning one point. In practice, multiple unlearning requests can arrive sequentially (sequential unlearning) and each unlearning request may require unlearning multiple points (batch unlearning). Below we demonstrate that our PNSGD unlearning naturally supports sequential and batch unlearning as well.
@@
Under the same assumptions as in Corollary 3.7. Assume the unlearning requests arrive sequentially such that our dataset changes from $\mathcal{D}=\mathcal{D}_{0}\rightarrow\mathcal{D}_{1}\rightarrow\ldots\rightarrow\mathcal{D}_{S}$, where $\mathcal{D}_{s},\mathcal{D}_{s+1}$ are adjacent. Let $y_{k}^{j,(s)}$ be the unlearned parameters for the $s^{th}$ unlearning request at $k^{th}$ unlearning epoch and $j^{th}$ iteration following Algorithm (1) on $\mathcal{D}_{s}$ and $y_{0}^{0,(s+1)}=y_{K_{s}}^{0,(s)}\sim\bar{\nu}_{\mathcal{D}_{s}|\mathcal{B}}$, where $y_{0}^{0,(0)}=x_{\infty}$ and $K_{s}$ is the unlearning steps for the $s^{th}$ unlearning request. For any $s\in[S]$, we have $W_{\infty}(\bar{\nu}_{\mathcal{D}_{s-1}|\mathcal{B}},\nu_{\mathcal{D}_{s}|\mathcal{B}})\leq Z_{\mathcal{B}}^{(s)},$ where $Z_{\mathcal{B}}^{(s+1)}=\min(c^{K_{s}n/b}Z_{\mathcal{B}}^{(s)}+Z_{\mathcal{B}},2R)$, $\;Z_{\mathcal{B}}^{(1)}=Z_{\mathcal{B}}$, $Z_{\mathcal{B}}$ and $c$ are described in Corollary 3.7.
Figure 2: Illustration of (a) sequential unlearning and (b) batch unlearning. The key idea is to establish an upper bound on the initial $W_{\infty}$ distance. (a) For sequential unlearning, the initial $W_{\infty}$ distance bound $Z_{\mathcal{B}}^{(s)}$ for each $s^{th}$ unlearning request can be derived with triangle inequality. (b) For batch unlearning, we analyze the case that two learning processes can differ in $S\geq 1$ points.
@@
By combining Corollary 3.7 and Theorem 3.11, we can establish the least unlearning iterations of each unlearning request $\{K_{s}\}_{s=1}^{S}$ to achieve $(\alpha,\varepsilon)$-RU simultaneously. Notably, our sequential unlearning bound is much better than the one in [11], especially when the number of unlearning requests is large. The key difference is that [11] have to leverage weak triangle inequality for Rényi divergence, which double the Rényi divergence order $\alpha$ for each sequential unlearning request. In contrast, since our analysis only requires tracking the initial $W_{\infty}$ distance, where the standard triangle inequality can be applied. As a result, our analysis can better handle the sequential unlearning case. We also demonstrate in Section 4 that the benefit offered by our results is significant in practice.
@@
Batch unlearning. We can extend Lemma 3.3 to the case that adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$ can differ in $S\geq 1$ points, which further leads to batch unlearning guarantee. We relegate the result in Appendix J.
@@
Figure 3: Main experiments, where the top and bottom rows are for MNIST and CIFAR10 respectively. (a) Compare to baseline for unlearning one point using limited $K$ unlearning epoch. For PNSGD, we use only $K=1$ unlearning epoch. For D2D, we allow it to use $K=1,5$ unlearning epochs. (b) Unlearning $100$ points sequentially versus baseline. For LU, since their unlearning complexity only stays in a reasonable range when combined with batch unlearning of size $S$ sufficiently large, we report such a result only. (c,d) Noise-accuracy-complexity trade-off of PNSGD for unlearning $100$ points sequentially with various mini-batch sizes $b$, where all methods achieve $(\epsilon,1/n)$-unlearning guarantee with $\epsilon=0.01$. We also report the required accumulated epochs for retraining for each $b$.
@@ 4 Experiments
Benchmark datasets. We consider binary logistic regression with $\ell_{2}$ regularization. We conduct experiments on MNIST [25] and CIFAR10 [26], which contain 11,982 and 10,000 training instances respectively. We follow the setting of [7, 11] to distinguish digits $3$ and $8$ for MNIST so that the problem is a binary classification. For the CIFAR10 dataset, we distinguish labels $3$ (cat) and $8$ (ship) and leverage the last layer of the public ResNet18 [27] embedding as the data features, which follows the setting of [7] with public feature extractor.
@@
Baseline methods. Our baseline methods include Delete-to-Descent (D2D) [9] and Langevin Unlearning (LU) [11], which are the state-of-the-art full-batch gradient-based approximate unlearning methods. Note that when our PNSGD unlearning chooses $b=n$ (i.e., full batch), the learning and unlearning iterations become PNGD which is identical to LU. Nevertheless, the corresponding privacy bound is still different as we leverage the analysis different from those based on Langevin dynamics in [11]. Hence, we still treat these two methods differently in our experiment. For D2D, we leverage Theorem 9 and 28 in [9] for privacy accounting depending on whether we allow D2D to have an internal non-private state. Note that allowing an internal non-private state provides a weaker notion of privacy guarantee [9] and both PNSGD and LU by default do not require it. We include those theorems for D2D and a detailed explanation of its possible non-privacy internal state in Appendix O. For LU, we leverage their Theorem 3.2, and 3.3 for privacy accounting [11], which are included in Appendix P.
@@
All experimental details can be found in Appendix N, including how to convert $(\alpha,\varepsilon)$-RU to the standard $(\epsilon,\delta)$-unlearning guarantee. Our code is publicly available [^3] https://github.com/Graph-COM/SGD_unlearning. We choose $\delta=1/n$ for each dataset and require all tested unlearning approaches to achieve $(\epsilon,\delta)$-unlearning with different $\epsilon$. We report test accuracy for all experiments as the utility metric. We set the learning iteration $T=10,20,50,1000$ to ensure PNSGD converges for mini-batch size $b=32,128,512,n$ respectively. All results are averaged over $100$ independent trials with standard deviation reported as shades in figures.
@@
Unlearning one data point with $K=1$ epoch. We first consider the setting of unlearning one data point using only one unlearning epoch (Figure 3a). For both LU and PNSGD, we use only $K=1$ unlearning epoch. Since D2D cannot achieve a privacy guarantee with only limited (i.e., less than $10$) unlearning epoch without a non-private internal state, we allow D2D to have it and set $K=1,5$ in this experiment. Even in this case, PNSGD still outperforms D2D in both utility and unlearning complexity. Compared to LU, our mini-batch setting either outperforms or is on par with it. Interestingly, we find that LU gives a better privacy bound compared to full-batch PNSGD ($b=n$) and thus achieves better utility under the same privacy constraint, see Appendix P for the detailed comparisons. Nevertheless, due to the use of weak triangle inequality in LU analysis, we will see that our PNSGD can outperform LU significantly for multiple unlearning requests.
@@
Unlearning multiple data points. Let us consider the case of multiple ($100$) unlearning requests (Figure 3b). We let all methods achieve the same $(1,1/n)$-unlearning guarantee for a fair comparison. We do not allow D2D to have an internal non-private state anymore in this experiment for a fair comparison. Since the privacy bound of LU only gives reasonable unlearning complexity with a limited number of sequential unlearning updates [11], we allow it to unlearn $S=10$ points at once. We observe that PNSGD requires roughly $10\%$ and $2\%$ of unlearning epochs compared to D2D and LU for $b=n$ and $b=128$ respectively, where all methods exhibit similar utility ($0.9$ and $0.98$ for MNIST and CIFAR10 respectively). It shows that PNSGD is much more efficient compared to D2D and LU. Notably, while both PNSGD with $b=n$ and LU (un)learn with PNGD iterations, the resulting privacy bound based on our PABI-based analysis is superior to the one pertaining to Langevin-dynamic-based analysis in [11]. See our discussion in Section 3.2 for the full details.
@@
Privacy-utility-complexity trade-off. We now investigate the inherent utility-complexity trade-off regarding noise standard deviation $\sigma$ and mini-batch size $b$ for PNSGD under the same privacy constraint, where we require all methods to achieve $(0.01,1/n)$-unlearning guarantee for $100$ sequential unlearning requests (Figure 3c and 3d). We can see that smaller $\sigma$ leads to a better utility, yet more unlearning epochs are needed for PNSGD to achieve $\epsilon=0.01$. On the other hand, smaller mini-batch size $b$ requires fewer unlearning epochs $K$ as shown in Figure 3d, since more unlearning iterations are performed per epoch. Nevertheless, we remark that choosing $b$ too small may lead to degradation of model utility or instability. Decreasing the mini-batch size $b$ from $32$ to $1$ reduces the average accuracy of training from scratch from $0.87$ to $0.64$ and $0.97$ to $0.81$ on MNIST and CIFAR10 respectively for $\sigma=0.03$. In practice, one should choose a moderate mini-batch size $b$ to ensure both good model utility and unlearning complexity. Finally, we also note that PNSGD achieves a similar utility while much better complexity compared to retraining from scratch, where PNSGD requires at most $1,5$ unlearning epochs per unlearning request for $b=32,512$ respectively.
5 Limitations and Conclusion
Limitation. Since our analysis is built on the works of [15, 16], we share the same limitation that the (strong) convexity assumption is required. It is an open problem on how to extend such analysis beyond convexity assumption as stated in [15, 16]. While we resolve this open problem for establishing DP properties of PNSGD in our recent work [28], it is still unclear whether the same success can be generalized to the unlearning problem. One interesting direction is to leverage the Langevin dynamic analysis [29] instead as in [11], which can deal with non-convex problems in theory yet we conjecture the resulting bounds can be loose, and more complicated.
Conclusion. We propose to leverage projected noisy stochastic gradient descent (PNSGD) for machine unlearning problem. We provide its unlearning guarantees as well as many other algorithmic benefits of PNSGD for unlearning under the convexity assumption. Our results are closely related to our new results on infinite Wasserstein distance tracking of the adjacent (un)learning processes, which is also leveraged in our concurrent work for studying DP-PageRank algorithms [30]. Extensive experiments show that our approach achieves a similar utility under the same privacy constraint while using $2\%$ and $10\%$ of the gradient computations compared with the state-of-the-art gradient-based approximate unlearning methods for mini-batch and full-batch settings, respectively.
Acknowledgments and Disclosure of Funding
@@ 5 Limitations and Conclusion
The authors thank Sinho Chewi, Wei-Ning Chen, and Ayush Sekhari for the helpful discussions. E. Chien, H. Wang and P. Li are supported by NSF awards OAC-2117997 and JPMC faculty award.
@@ Appendix A Additional related works
Differential privacy of noisy gradient methods. DP-SGD [13] is arguably the most popular method for ensuring a DP guarantee for machine learning models. Since it leverages the DP composition theorem and thus the privacy loss will diverge for infinite training epochs. Recently, researchers have found that if we only release the last step of the trained model, then we can do much better than applying the composition theorem. A pioneer work [31] studied the DP properties of Langevin Monte Carlo methods. Yet, they do not propose to use noisy GD for general machine learning problems. A recent line of work [22, 32] shows that PNSGD training can not only provide DP guarantees, but also the privacy loss is at most a finite value even if we train with an infinite number of iterations. The main analysis therein is based on the analysis of Langevin Monte Carlo [33, 29]. In the meanwhile, [15] also provided the DP guarantees for PNSGD training but with analysis based on privacy amplification by iteration [20]. None of these works study how PNSGD can also be leveraged for machine unlearning.
Appendix B Limitations
Since our analysis is built on the works of [15, 16], we share the same limitation that the (strong) convexity assumption is required. It is an open problem on how to extend such analysis beyond convexity assumption as stated in [15, 16]. While we resolve this open problem for the DP properties of PNSGD in our recent work [28], it is still unclear whether the same success can be generalized to the unlearning problem. One interesting direction is to leverage the Langevin dynamic analysis [29] instead as in [11], which can deal with non-convex problems in theory yet we conjecture the resulting bounds can be loose, and more complicated.
Appendix C Broader Impact
Our work study the theoretical unlearning guarantees of projected stochastic noisy gradient descent algorithm for convex problems. We believe our work is a foundational research and does not have a direct path to any negative applications.
Appendix D Standard definitions
Let $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ be a mapping. We define smoothness, Lipschitzness, and strong convexity as follows:
@@ Appendix D Standard definitions
Appendix E Existence of limiting distribution
@@
Suppose that the closed convex set $\mathcal{C}\subset\mathbb{R}^{d}$ is bounded with $\mathcal{C}$ having a positive Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. The Markov chain $\{x_{t}:=x_{t}^{0}\}$ in Algorithm 1 for any fixed mini-batch sequence =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=1}^{n/b-1}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on the Borel $\sigma$-algebra of $\mathcal{C}$. Furthermore, for any $x\in\mathcal{C}$, the distribution of $x_{t}$ conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$.
The proof is almost identical to the proof of Theorem 3.1 in [11] and we include it for completeness. We start by proving that the process $\{x_{t}:=x_{t}^{0}\}$ admits a unique invariant measure (Proposition E.1) and then show that the process converges to such measure which is in fact a probability measure (Theorem E.2). Combining these two results completes the proof of Theorem 3.1.
Proposition E.1.
Suppose that the closed convex set $\mathcal{C}\subset\mathbb{R}^{d}$ is bounded with $\textrm{Leb}(\mathcal{C})>0$ where Leb denotes the Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. Then the Markov chain $\{x_{t}:=x_{t}^{0}\}$ defined in Algorithm 1 for any fixed mini-batch sequence =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=1}^{n/b-1}$ admits a unique invariant measure (up to constant multiples) on $\mathcal{B}(\mathcal{C})$ that is the Borel $\sigma$-algebra of $\mathcal{C}$.
@@
This proposition is a direct application of results from [23]. According to Proposition 10.4.2 in [23], it suffices to verify that $\{x_{t}\}$ is recurrent and strongly aperiodic.
@@
    Recurrency. Thanks to the Gaussian noise $W_{t}$, $\{x_{t}\}$ is Leb-irreducible, i.e., it holds for any $x\in\mathcal{C}$ and any $A\in\mathcal{B}(\mathcal{C})$ with $\textrm{Leb}(A)>0$ that
@@
    where $\tau_{A}=\inf\{t\geq 0:x_{t}\in A\}$ is the stopping time. Therefore, there exists a Borel probability measure $\psi$ such that that $\{x_{t}\}$ is $\psi$-irreducible and $\psi$ is maximal in the sense of Proposition 4.2.2 in [23]. Consider any $A\in\mathcal{B}(\mathcal{C})$ with $\psi(A)>0$. Since $\{x_{t}\}$ is $\psi$-irreducible, one has $L(x,A)=\mathbb{P}(\tau_{A}<+\infty\ |\ x_{0}=x)>0$ for all $x\in\mathcal{C}$. This implies that there exists $T\geq 0$, $\delta>0$, and $B\in\mathcal{B}(\mathcal{C})$ with $\textrm{Leb}(B)>0$, such that $\mathbb{P}(x_{T}\in A\ |\ x_{0}=x)\geq\delta,\ \forall~{}x\in B$. Therefore, one can conclude for any $x\in\mathcal{C}$ that
@@
    $\geq$ $\sum_{t=1}^{\infty}\delta\cdot\inf_{y\in\mathcal{C}}\mathbb{P}(x_{t}\in B\ |\ x_{t-1}=y)$
@@
    where we used the fact that $\inf_{y\in\mathcal{C}}\mathbb{P}(x_{t}\in B\ |\ x_{t-1}=y)=\inf_{y\in\mathcal{C}}\mathbb{P}(x_{1}\in B\ |\ x_{0}=y)>0$ that is implies by $\textrm{Leb}(B)>0$ and the boundedness of $\mathcal{C}$ and $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C};\mathbf{d}_{i})$. Let us remark that we actually have compact $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C};\mathbf{d}_{i})$ since $\mathcal{C}$ is compact and $\nabla f(\cdot;\mathbf{d}_{i})$ is continuous. The arguments above verify that $\{x_{t}\}$ is recurrent (see Section 8.2.3 in [23] for definition).